jwaskiewicz:

rewrite ^3 sqrt 27x-81-5 to make it easy to graph a translation

geerky42:

Is it \(\large \sqrt[3]{27x-81}-5\)? If so then you can factor out 27 inside the root: \[\large \sqrt[3]{27(x-3)}-5\\~\\\large\sqrt[3]{3^3(x-3)}-5\\~\large3\sqrt[3]{x-3}-5\] Now it's easier to graph it. Take \(f(x)=\sqrt x\) then our function would be \(3f(x-3)-5\).

jwaskiewicz:

thanks so much